1.What is equilibrium molarity?Plz xplain a bit?


2.What is equivalent mass?How can we get the eqv mass of any species 4m a rxn?

Grade:Upto college level

1 Answers

Ramesh V
14 years ago

(1) I'll explain through examples :

Calculate the equilibrium molarities of H2, I2, and HI

                              H2(g) + I2(g) <--> 2 HI(g)
Initial..................... 0.05.... 0.06 ...........0
Change...................- x ......- x..............+x
Equilibrium.......... 0.05 - x.. 0.06-x...........x

Equilibrium constant, Kc = [HI]2 / [H2] [I2] = 58.6
......................................… = [x]2 / [0.05-x] [0.06-x] = 58.6

Solve for x to get the change in the number of moles, and subsequently the number of moles of each compound in equilibrium

Molarity of H2 = [0.05-x] / 1.00L
Molarity of I2 = [0.06-x] / 1.00L
Molarity of HI = x / 1.00L

(2)

Equivalent weight is the atomic weight of an element or radical divided by the valence it assumes in a chemical compound.

5H2SO4+8NaI=4Na2SO4+H2S+4H2O+4I2

By definition one equivalent (or equivalent weight) of a substance is the amount of that substance which supplies or consumes one mole of reactive species. In acid-base chemistry the reactive species is the hydrogen ion (H+) while in oxidation-reduction chemistry the reactive species is the electron.

Since this is an acid-base reaction, we need to follow the rules in getting equivalent weights appropriate for this kind of reaction. From the equation, we can conclude that there are 4 reactive hydrogen ions based from the produced H2O.

EW = Molar Weight (MW) / (No. of Equivalence / No. of Moles)
EW = (98.0791 g/mol) / (4 eq / 5 mol) = 122.6 g/eq

Hope that helps! :)
--
regards
Ramesh